Boeing Corporation has just issued a callable (at par) three-year, 5% coupon bond with semiannual coupon payments. The bond can be called at par in two years or anytime thereafter on a coupon payment date. It has a price of $99.

  1. a.         What is the bond’s yield to maturity?
  2. b.         What is its yield to call?
  3. c.         What is its yield to worst?
Answer #1

a]

YTM is calculated using RATE function in Excel :

nper = 3 * 2 (3 years remaining until maturity with 2 semiannual coupon payments each year)

pmt = 100 * 5% / 2 (semiannual coupon payment = face value * coupon rate / 2)

pv = -99 (Current price of bond. This is entered with a negative sign because it is a cash outflow to buy the bond today).

fv = 100 (face value of bond receivable at maturity).

RATE is calculated to be 2.68%. This is the semiannual YTM. To get annual YTM, we multiply by 2.

Annual YTM is 5.37%

b]

YTC is calculated using RATE function in Excel :

nper = 2 * 2 (2 years remaining until first call with 2 semiannual coupon payments each year)

pmt = 100 * 5% / 2 (semiannual coupon payment = face value * coupon rate / 2)

pv = -99 (Current price of bond. This is entered with a negative sign because it is a cash outflow to buy the bond today).

fv = 100 (call price of bond receivable at first call date).

RATE is calculated to be 2.77%. This is the semiannual YTC. To get annual YTC, we multiply by 2.

Annual YTC is 5.54%

c]

Yield to worst is the lower of YTM and YTC.

Yield to worst is 5.37%